1)     Assume the logarithmic transformation f a utility function, for the consumption of two commodities is given by

ln U = ln4 + 0.5ln X + 0.25lnY

(a) if the price of X is GHS2.50 and that of Y is GHS4.00, calculate the optimal combination for an income of GHS50.00.

b)  Determine and interpret the value of the Lagrange multiplier.

 


1
Expert's answer
2021-07-27T16:17:02-0400

"log U =log 4+0.5 log X+0.25 log Y"

"50=2.50X+4Y"

"L=log 4+ 0.5log X +0.25 log Y+ \lambda [50-2.50X-4Y]"

"\frac {\delta L}{\delta X}=\frac {0.5}{X}-\lambda (2.50)=0"


"\frac {\delta L}{\delta Y}=\frac {0.25}{Y}-\lambda(4)=0"


"\frac {0.5}{2.50X}=\frac {0.25}{4Y}"


"\frac {0.2}{X}=\frac {0.0625}{Y}"


"\frac {X}{Y}= \frac {0.2}{0.0625}"


"X= 3.2Y"

"50=2.50X+4Y"

"50=2.50(3.2Y)+4Y"

"50=12Y"

"Y=4.16"

"X=13.33"